Two vertical poles of lengths 6 feet and 8 feet stand 10 feet apart. a cable reaches from the top of one pole to some peoint of the ground between the poles then to the top of the other pole. where should this point be located to use 18 feet of cable?
could you explain how to do this for me?

Answer by stanbon(75887)

Draw the picture.
The point where the wire touches the ground divides
the base into two parts: x and 10-x
---------------------------------------
You have two right triangles and the wire forms the hypotenuse of
both of those triangles.
--------------------------------
Hypotenuse 1 = sqrt(6^2 + (10-x)^2)
Hypotenuse 2 = sqrt(8^2 + x^2)
------------------------------------
Equation:
sqrt(36 + (10-x)^2) + sqrt(64 + x^2) = 18
---
Solve for "x" to find the point where the wire should touch the ground.
===================================
Cheers,
Stan H.